OpenStudy (anonymous):

A piece of circular bread has a diameter of 11 in. What is the approximate area of this piece of bread? Use 3.14 to approximate pi and round your answer to the nearest hundredth. ___ in2

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

*WILL GIVE MEDAL AND FAN*

OpenStudy (anonymous):

Area of a circle: \[A = \pi r^2 \approx 3.14 \cdot r^2\] \[r = d/2\] Use these equations.

OpenStudy (anonymous):

I got 17.27. Is this my final answer?

OpenStudy (anonymous):

That doesn't sound right.

OpenStudy (bookworm14):

so do 11 ^2 = 11*11 then mulitply that *3.14

OpenStudy (bookworm14):

@sonic400539

OpenStudy (anonymous):

@Bookworm14 That's not right. 11 is the diameter, not the radius. The radius is half the diameter, or 5.5.

OpenStudy (bookworm14):

oh, woops i read the question to fast! x( @adotm so you would do 11 * 3.14 = 34.54 @sonic400539